OpenStudy (anonymous):

The elements in this circuit have the following values: R1=100Ω, R2=100Ω, R3=200Ω, V1=5V, I1=0.01A and I2=0.001A . please tell the answers for question 1(midterm).and also how to solve them

OpenStudy (anonymous):

1) (R*b)/(1+R*b)*vIN 2) (VS*(1+R*b)-vIN)/(1+R*b) current through R1 =current through R2 =(V1-R3(I1-I2))/(R1+R2+R3) current through R3 =(V1-(R1R2)(I1-I2))/(R1+R2+R3) 1. Va=(current through R2)*R2+Vb 2. Vb=(current through R3 )*R3 3. power in R1= current through R1*current through R1*R1 4. power in R2=current through R2*current through R2*R2 5. power in R3=current through R3*current through R2*R3 6. power in V1=V1*current through R1 7. power in i1=Vb*current through R3 8. power in i2=-Vb*current through R3

OpenStudy (anonymous):

thanks ali... but vb and va aren't comming correct.., can you check this...,

OpenStudy (anonymous):

how to get va and vb